12,−6,−7,−9,4,0 Determine the mean of the given data.

12,−6,−7,−9,4,0

Determine the mean of the given data.
Answer #1

12, -6, -7, -9, 410 Mean = (12-6-7-9+4+0)/6 = -6/6 = -1 ans Mean = ]